OpenStudy (anonymous):

Give the solution to the system that is represented by 1 0 2 5 0 1 1 3 0 0 0 0 i keep getting x=2z+5, z=-y+3 but don't know where to go from there. Thanks

OpenStudy (anonymous):

Don´t understand the signs you put. Can explain, please?

OpenStudy (anonymous):

so my teacher gave me a matrices and when it is a infinite # of solutions he wants us to write it out like (heres an example) -10Z+5, 5z-1,z)

OpenStudy (anonymous):

Well, the columns are x, y, z and the independet term, aren´t they? So I´ll write the next: x+ 2z = 5; y+z = 3. And the answer would be x in terms of y and z. Do you think so?

OpenStudy (anonymous):

im not sure. Im reviewing for my exam and have been stuck on this one for a while and cannot find an example or anything. Thank you for your help though

OpenStudy (anonymous):

jkd is htis an augmented matrice

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x+2z=5 y+z=3 or \[C_1+2C_3=5\] \[C_2+C_3=3\] you can see that both hae c_3 in each ... let us create parameters where t is our independent variable t.

OpenStudy (anonymous):

if \[c_3=t\] \[c_1=5-2t\] \[c_2=3-t\]

OpenStudy (anonymous):

now t can be any number that is why there is infinite solutions

OpenStudy (anonymous):

do you need this in vector forms?

OpenStudy (anonymous):

your \[y=(5,3,0)^{T}+t(-2,-1,1)^T\]

OpenStudy (anonymous):

thank you for your help i think i got it!